# Proprioception Enhances Vision Language Model in Generating Captions and Subtask Segmentations for Robot Task

### v1

@@ I INTRODUCTION
In recent years, Large Language Models (LLMs), Vision Language Models (VLMs), and other foundation models have been applied to planning and recognition tasks in robotics, achieving notable results [CLIPort][Ahn2022a][kirillov2023segany]. However, it remains unclear whether these models truly understand the dynamics of robot motion. Most foundation models do not include the robot’s proprioception, such as sensorimotor trajectories, in their training datasets. Robot motion information associated with the body is extremely important for building a world model that includes physical dynamics. Due to this lack of that input, previous studies on VLMs and LLMs handle motion trajectory outputs in predefined API formats [Liang2022a], which are often described as superficial [alice2024][mollo2023][harnad2025]. Therefore, investigating the robot motion understanding capabilities of VLMs is a crucial research topic for advancing robot foundation models [Driess2023a][pi0].
One way to measure a VLM’s understanding of the robot’s motion is through video captioning [yoshino]. The skills necessary for this task can be categorized into two types:
@@
The first skill is the ability to describe a series of robot motions in natural language when provided with task-related images or videos. The language-motion pair data generated through captioning is crucial for integrated learning that connects robot actions with linguistic instructions [jang2021bc][Toyoda2022][Shi2024][Suzuki2024].
Moreover, for a robot to perform tasks effectively in real environments, it requires specific, executable motion instructions rather than mere descriptions of its actions. The second skill is the ability to divide a series of robot tasks into multiple subtasks. Previous studies have proposed methods in which subtask divisions are either predefined by humans [Fujii2022sii] or learned through training [beal][nakajo]. If VLMs can extract subtasks from complex, long-term tasks in a zero-shot manner, more efficient robot learning is likely to be achieved.
@@
In this study, we propose a method for robot motion captioning and subtask segmentation using a VLM. We believe that proprioception is important for understanding robot actions, and we take the approach of incorporating it into the VLM prompts. Through the two captioning tasks, we aim to verify whether the VLM can comprehend robot motion and generate outputs that account for the robot’s trajectory (Fig. 1). The proposed method builds on a video caption generation approach [amirian], generating “scene” captions from evenly spaced image captions and summarizing them into a complete task description. We also segment a series of motions into subtasks by comparing the text similarity between image captions. Additionally, the method enhances caption accuracy by incorporating low-level robot trajectory data (joint and end-effector states paired with images) into the prompt that is input to the VLM. In the experiment, we validate the effectiveness of the proposed method by applying it to several robot tasks recorded through human operation in a simulator.
@@ II-A Video Captioning
In image processing, video captioning is widely studied as a means of linking time-series data with natural language [amirian][Vid2Seq]. The primary goal in this field is to generate titles and summaries for movies and videos. Amirian et al. [amirian] generated sub-captions for representative image frames and then summarized them to produce a title and abstract for the video. Yang et al. [Vid2Seq] enhanced a language model with time tokens to enable seamless caption prediction. They defined the end of a sentence in transcribed audio as a pseudo-event boundary to align captions with corresponding video segments. Although these studies achieve high-accuracy caption generation, most previous work did not focus on robot motions. Additionally, the input data typically lacked key information, such as the motion trajectory.
@@ II-B Segmentation of Robot Motion
In contrast, many studies have focused on segmenting robot motions using machine learning [yoshino][beal][nakajo][nagano]. Beal et al. [beal] proposed a motion segmentation method based on a Hidden Markov Model (HMM) that accounts for infinite time states. Nagano et al. [nagano] extended the HMM by integrating a variational autoencoder, enabling motion segmentation that considers the duration of each segment. Yoshino et al. [yoshino] introduced a method that embeds actuator and camera transitions into a model through unsupervised learning, clustering motion classes using the $k$-nearest neighbor method. Nakajo et al. [nakajo] incorporated parametric bias (PB) into an imitation learning framework, proposing an approximate subtask segmentation method by grouping PBs with similar values. While these studies rely on the learning trajectory of the robot’s task experience, they do not address the performance evaluation of trained VLMs, which is the focus of this study. The robot tasks segmented in the studies mentioned above can be incorporated into an imitation learning framework as individual training data. It has also been suggested that subtasks could be embedded into a model in a manner that allows them to be switched and combined [Kase2018icra]. Thus, achieving automatic subtask segmentation has the potential to significantly enhance the performance of robot learning.
In recent years, there has been active progress in developing the Vision Language Action Model (VLA [vla], an extension of the VLM. The VLA aims to create a foundation model that incorporates physical dynamics by fine-tuning the VLM to predict robot motion information [rtx]. Additionally, the world foundation model leverages a VLM trained on over 20 million hours of video data, enabling video output that reflects physical dynamics [cosmos]. In [alter3], the input prompt for the LLM is modified to include joint angle information, allowing it to directly output the robot’s trajectory. This suggests that foundation models may have the potential to predict robot dynamics through prompt engineering. In this study, we use the zero-shot inference capabilities of VLM to perform captioning of robot motions. Our contributions can be summarized as follows:
@@ III PROPOSED METHOD
An overview of the proposed method is shown in Fig. 2. Our proposed method is based on a previous study of video captioning [amirian]. We generate a final caption of the entire robot task by performing step-by-step caption generation (Section III-A). The proposed method enhances caption accuracy by incorporating low-level robot trajectory data into the prompt that is input to the VLM. Then, we automatically divide the robot task by measuring the similarity of the image captions generated during the captioning process (Section III-B).
@@ IV EXPERIMENTS
We conducted experiments to generate captions for some robot tasks recorded on a simulator and verify the effectiveness of the proposed method. We targeted two tasks implemented in the Robosuite simulator [robosuite2020]: Door opening and Bin pick-placing for cereal (Fig. 3). As robot motion data, we recorded 22 sequences operated by a human using a 3D mouse. 10 sequences are for Door opening and 12 sequences are for Bin pick-placing (3 sequences for each bin position). The motion data included RGB images of $\times256$ pixels, 8 degree-of-freedom joint angles including gripper opening and closing, and the end-effector’s posture. The camera was positioned so that the images could capture the entire robot task. The sequence length for each task was 580 and 539 steps on average for Door opening and Bin pick-placing, respectively.
In the caption generation task, we evaluated whether the final captions accurately described key elements of each task, including the object being manipulated (e.g., doorknob, cereal), the action (e.g., opening, pick-and-place), and trajectory information (e.g., motion direction, bin position). If all these elements were correctly captured, we determined that the VLM successfully generated captions containing detailed motion trajectory information. In the subtask segmentation task, we converted the image captions generated by the VLM into 768-dimensional embeddings using the stsb-xlm-r-multilingual model from the sentence-transformers library and calculated the cosine similarity between consecutive image captions. We examined how the subtask segmentation behavior of the VLM varied based on the similarity threshold between captions. For our experiments, we used ChatGPT-4V [ChatGPT] as the VLM, setting the temperature to zero for all tests. ChatGPT is one of the most widely used LLM/VLM models in academic research [Vemprala2023]. The version employed in our study was gpt-4v-preview.
@@ VI CONCLUSION
In this study, we proposed a method for automatic captioning of robot motions using a VLM that incorporates motion information into prompts and explored whether the VLM understands robot motions. In the proposed method, the scene captions were generated from the image captions and trajectory data of a series of robot motions, and the entire task caption was generated by summarizing them. We also divided the series of motions into subtasks by comparing the similarity between text embedding of image captions. Experiments using two robot tasks on a simulator demonstrated the effectiveness of the proposed method. Results confirmed that the accuracy of generated captions and subtask segmentation improved when motion information (joint angles and end-effector state) was included in the VLM input. The findings also suggested that while the VLM has a superficial understanding of robot motion, it struggles to predict fine changes over time. In future work, we plan to extend the proposed method by incorporating it into imitation learning frameworks [Suzuki2023] and LLM-based motion planning [Hori2023].

### v2

@@ I INTRODUCTION
In recent years, Large Language Models (LLMs), Vision Language Models (VLMs), and other foundation models have been applied to planning and recognition tasks in robotics, achieving notable results [1][2][3]. However, it remains unclear whether these models truly understand the dynamics of robot motion. Most foundation models do not include the robot’s proprioception, such as sensorimotor trajectories, in their training datasets. Robot motion information associated with the body is extremely important for building a world model that includes physical dynamics. Due to this lack of that input, previous studies on VLMs and LLMs handle motion trajectory outputs in predefined API formats [4], which are often described as superficial [5][6][7]. Therefore, investigating the robot motion understanding capabilities of VLMs is a crucial research topic for advancing robot foundation models [8][9][10][11][12].
One way to measure a VLM’s understanding of the robot’s motion is through video captioning [13]. The skills necessary for this task can be categorized into two types:
@@
The first skill is the ability to describe a series of robot motions in natural language when provided with task-related images or videos. The language-motion pair data generated through captioning is crucial for integrated learning that connects robot actions with linguistic instructions [14][15][16][17].
Moreover, for a robot to perform tasks effectively in real environments, it requires specific, executable motion instructions rather than mere descriptions of its actions. The second skill is the ability to divide a series of robot tasks into multiple subtasks. Previous studies have proposed methods in which subtask divisions are either predefined by humans [18] or learned through training [19][20]. If VLMs can extract subtasks from complex and long-horizon tasks in a zero-shot manner, more efficient robot learning is likely to be achieved.
@@
In this study, we propose a method for robot motion captioning and subtask segmentation using a VLM. We believe that proprioception is important for understanding robot actions, and we take the approach of incorporating it into the VLM prompts. Through the two captioning tasks, we aim to verify whether the VLM can comprehend robot motion and generate outputs that account for the robot’s trajectory (Fig. 1). The proposed method builds on a video caption generation approach [21], generating “scene” captions from evenly spaced image captions and summarizing them into a complete task description. We also segment a series of motions into subtasks by comparing the text similarity between image captions. Additionally, the method enhances caption accuracy by incorporating low-level robot trajectory data (joint and end-effector states paired with images) into the prompt that is input to the VLM. In the experiment, we validate the effectiveness of the proposed method by applying it to several robot tasks recorded through human operation in a simulator.
@@ II-A Video Captioning
In image processing, video captioning is widely studied as a means of linking time-series data with natural language [21][22]. The primary goal in this field is to generate titles and summaries for movies and videos. Amirian et al. [21] generated sub-captions for representative image frames and then summarized them to produce a title and abstract for the video. Yang et al. [22] enhanced a language model with time tokens to enable seamless caption prediction. They defined the end of a sentence in transcribed audio as a pseudo-event boundary to align captions with corresponding video segments. Although these studies achieve high-accuracy caption generation, most previous work did not focus on robot motions. Additionally, the input data typically lacked key information, such as the motion trajectory.
@@ II-B Segmentation of Robot Motion
In contrast, many studies have focused on segmenting robot motions using machine learning [13][19][20][23]. Beal et al. [19] proposed a motion segmentation method based on a Hidden Markov Model (HMM) that accounts for infinite time states. Nagano et al. [23] extended the HMM by integrating a variational autoencoder, enabling motion segmentation that considers the duration of each segment. Yoshino et al. [13] introduced a method that embeds actuator and camera transitions into a model through unsupervised learning, clustering motion classes using the $k$-nearest neighbor method. Nakajo et al. [20] incorporated parametric bias (PB) into an imitation learning framework, proposing an approximate subtask segmentation method by grouping PBs with similar values. While these studies rely on the learning trajectory of the robot’s task experience, they do not address the performance evaluation of trained VLMs, which is the focus of this study. The robot tasks segmented in the studies mentioned above can be incorporated into an imitation learning framework as individual training data. It has also been suggested that subtasks could be embedded into a model in a manner that allows them to be switched and combined [24]. Thus, achieving automatic subtask segmentation has the potential to significantly enhance the performance of robot learning.
II-C VLA and World Model
In recent years, there has been active progress in developing the Vision Language Action Model (VLA [8][9][10][11][12]), The VLA aims to create a foundation model that incorporates physical dynamics by fine-tuning the VLM to predict robot motion information. Recent advances in VLA have shown that pre-training a VLM to predict latent actions learned from video data can improve performance [25]. However, this does not measure the performance of the original VLM.
Additionally, the world foundation model leverages a VLM trained on over 20 million hours of video data, enabling video output that reflects physical dynamics [26]. Some studies have incorporated world models into VLA to connect videos of human tasks with robot behavior [27][28]. As VLM’s ability to understand robot behavior improves, it is expected that the prediction accuracy of these studies will also improve.
And, in [29], the input prompt for the LLM is modified to include joint angle information, allowing it to directly output the robot’s trajectory. This suggests that foundation models may have the potential to predict robot dynamics through prompt engineering. In this study, we use the zero-shot inference capabilities of VLM to perform captioning of robot motions. Our contributions can be summarized as follows:
@@ III PROPOSED METHOD
An overview of the proposed method is shown in Fig. 2. Our proposed method is based on a previous study of video captioning [21]. We generate a final caption of the entire robot task by performing step-by-step caption generation (Section III-A). The proposed method enhances caption accuracy by incorporating low-level robot trajectory data into the prompt that is input to the VLM. Then, we automatically divide the robot task by measuring the similarity of the image captions generated during the captioning process (Section III-B).
@@ IV EXPERIMENTS
We conducted experiments to generate captions for some robot tasks recorded on a simulator and verify the effectiveness of the proposed method. We targeted two tasks implemented in the Robosuite simulator [30]: Door opening and Bin pick-placing for cereal (Fig. 3). As robot motion data, we recorded 22 sequences operated by a human using a 3D mouse. 10 sequences are for Door opening and 12 sequences are for Bin pick-placing (3 sequences for each bin position). The motion data included RGB images of $\times256$ pixels, 8 degree-of-freedom joint angles including gripper opening and closing, and the end-effector’s posture. The camera was positioned so that the images could capture the entire robot task. The sequence length for each task was 580 and 539 steps on average for Door opening and Bin pick-placing, respectively.
In the caption generation task, we evaluated whether the final captions accurately described key elements of each task, including the object being manipulated (e.g., doorknob, cereal), the action (e.g., opening, pick-and-place), and trajectory information (e.g., motion direction, bin position). If all these elements were correctly captured, we determined that the VLM successfully generated captions containing detailed motion trajectory information. In the subtask segmentation task, we converted the image captions generated by the VLM into 768-dimensional embeddings using the stsb-xlm-r-multilingual model from the sentence-transformers library and calculated the cosine similarity between consecutive image captions. We examined how the subtask segmentation behavior of the VLM varied based on the similarity threshold between captions. For our experiments, we used ChatGPT-4V [31] as the VLM, setting the temperature to zero for all tests. ChatGPT is one of the most widely used LLM/VLM models in academic research [32]. The version employed in our study was gpt-4v-preview.
@@ VI CONCLUSION
In this study, we proposed a method for automatic captioning of robot motions using a VLM that incorporates motion information into prompts and explored whether the VLM understands robot motions. In the proposed method, the scene captions were generated from the image captions and trajectory data of a series of robot motions, and the entire task caption was generated by summarizing them. We also divided the series of motions into subtasks by comparing the similarity between text embedding of image captions. Experiments using two robot tasks on a simulator demonstrated the effectiveness of the proposed method. Results confirmed that the accuracy of generated captions and subtask segmentation improved when motion information (joint angles and end-effector state) was included in the VLM input. The findings also suggested that while the VLM has a superficial understanding of robot motion, it struggles to predict fine changes over time. In future work, we plan to extend the proposed method by incorporating it into imitation learning frameworks [33] and LLM-based motion planning [34].